OpenStudy (anonymous):

Use the quadratic formula to solve 16x2 + 96x + 128 = 0

OpenStudy (anonymous):

deja vu

OpenStudy (anonymous):

\[16x^2 + 96x + 128 = 0\\ \\16(x^2+6x+8)=0 \\x^2+6x+8=0\]

OpenStudy (uri):

@dangelosmith http://blogs.discovermagazine.com/loom/files/2008/07/quadratic.jpg Here

OpenStudy (anonymous):

then factor as \((x+2)(x+4)=0\) fini

OpenStudy (uri):

a:16 b:96 c =128

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wait i cant find the answer

OpenStudy (whpalmer4):

if the teacher asks you to solve via the quadratic formula, then you should probably use that rather than factoring. But, as @satellite73 has kindly shown you all but the very final step for the factoring approach, and you don't see how to find the answer, I'll tell you. if you've factored your equation into something like \((x-1)(x+3) = 0\), the only step remaining is to recognize that for that to be true, \(x-1 = 0\) or \(x+3=0\). Set each factor equal to zero and solve. For my example (NOT the same answers as your problem), \[x-1=0\]\[x=1\]and\[x+3=0\]\[x=-3\]so \(x = 1, x = -3\) would be the solutions. Do the same with the factors @satellite73 provided to get the answers for this problem. Or, use the quadratic formula as the problem requests. I described the procedure in another of your posts.